OpenStudy (anonymous):

MEDAL AND FAN FOR BEST ANSWER: [8.07] What are the factors of x^2 − 81?

OpenStudy (anonymous):

@ajeetkumarseo @iambatman

OpenStudy (anonymous):

@Chad123 please help

OpenStudy (anonymous):

(X-9)(x+9)

OpenStudy (anonymous):

Thank You @Chad123 !!

OpenStudy (anonymous):

and @iambatman thank u!!

OpenStudy (anonymous):

\[(x^2-a^2) = (a+a)(x-a)\]

OpenStudy (anonymous):

that should be x+a not a + a

OpenStudy (skullpatrol):

$$\Huge x^2 − 81$$$$\Huge =x^2-9^2$$ $$\Huge a^2 − b^2=(a+b)(a-b)$$